A car A is travelling on straight level road with a uniform speed of 60 kilometre per hour it is followed by another B which is moving with a speed of 70 kilometre per hour when the distance between them is 2.5 kilometres that can be is given its deceleration of 20 kilometre per hour after the what distance and time will B catch up with A

A car A  is travelling on straight level road with a uniform speed of 60 kilometre per hour it is followed by another B which is moving with a speed of 70 kilometre per hour when the distance between them is 2.5 kilometres that can be is given its deceleration of 20 kilometre per hour after the what distance and time will B catch up with A

Grade:9

1 Answers

Arun
5 years ago
Dear student
 
 
s2 = ut + ½*a*t2 = 70t + ½*(-20)*t2 = 70t -10t2
s1 = 60t
s2 – s1 = 2.5
70t – 10t2 -60t = 2.5
10t2 -10t + 2.5 = 0
so t = ½ h
putting this in s2
s2 = 32.5km
Hope it clears
 
Regards
Arun